OpenStudy (anonymous):

2x-9(under radical)=1/3

OpenStudy (anonymous):

do you mean 2x - 9 = 1/3

OpenStudy (anonymous):

\[\sqrt{2x-9}=1/3\]

OpenStudy (anonymous):

Ok that makes more sense

Parth (parthkohli):

Square both sides to eliminate the radical and then solve for \(x\) :)

OpenStudy (mimi_x3):

\(\sqrt{2x-9}= \frac{1}{3}\) \(2x-9 = \frac{1}{3}^{2}\)

OpenStudy (anonymous):

What dont you understand about this question

OpenStudy (anonymous):

just making sure i'm doing it correctly.

OpenStudy (anonymous):

Ok. Just square both sides, this will give 2x - 9 = (1/9) now multiply both sides by 9, this will give 18x - 81 = 1 Add 81 to both sides 18x = 82 x = 82/18